# Overcoming Knowledge Barriers: Online Imitation Learning from Observation with Pretrained World Models

### v1

@@ Abstract
Incorporating the successful paradigm of pretraining and finetuning from Computer Vision and Natural Language Processing into decision-making has become increasingly popular in recent years. In this paper, we study Imitation Learning from Observation with pretrained models and find existing approaches such as BCO and AIME face knowledge barriers, specifically the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB), greatly limiting their performance. The EKB arises when pretrained models lack knowledge about unseen observations, leading to errors in action inference. The DKB results from policies trained on limited demonstrations, hindering adaptability to diverse scenarios. We thoroughly analyse the underlying mechanism of these barriers and propose AIME-v2 upon AIME as a solution. AIME-v2 uses online interactions with data-driven regulariser to alleviate the EKB and mitigates the DKB by introducing a surrogate reward function to enhance policy training. Experimental results on tasks from the DeepMind Control Suite and Meta-World benchmarks demonstrate the effectiveness of these modifications in improving both sample-efficiency and converged performance. The study contributes valuable insights into resolving knowledge barriers for enhanced decision-making in pretraining-based approaches. Code will be available at https://github.com/argmax-ai/aime-v2.
@@ 1 Introduction
In recent years, we are going through a paradigm shift from learning from scratch to pretraining and finetuning in Computer Vision (CV) [25, 47, 26] and Natural Language Processing (NLP) [12, 46, 43, 59, 60] fields due to the increasing availability of foundation models [6] and ever growing datasets. However, how to adapt this new paradigm into decision making is still unclear, i.e. what type of models we need to pretrain and how these models can be adapted to solve downstream tasks. Recent works [65, 11, 52, 48, 23] showcase that one can pretrain a shared latent space world model, enabling successful and efficient transfer to new tasks with either reinforcement learning [52, 48, 23] or Imitation Learning from Observation (ILfO) [65, 11]. Arguably, ILfO [57, 58, 3, 65, 11, 37], especially from videos [3, 65, 37, 11], is a more promising approach in this new paradigm since it does not require a handcrafted reward function which is hard to define for many real-world tasks.
@@
However for ILfO with pretrained models, we identify two barriers, which limit the effectiveness of employing pretrained models, the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB). The EKB describes the shortcomings of a pretrained model when confronted with novel observations and actions beyond its training experience. The DKB describes the known problem of generalisation from a limited number of expert demonstrations in imitation learning [27]. We find current approaches like BCO [57] and AIME [65] typically suffer from these two knowledge barriers. First, these algorithms depend on the pretrained model to successfully infer missing actions from observation sequences. Thus, when the model has not seen a specific observation before, it may not contain sufficient knowledge about the embodiment to infer the correct action. Second, policy optimisation is only guided by a limited demonstrations, which may not provide sufficient coverage of the space and lead to a brittle policy working well in some scenarios but not others.
@@
Figure 1: Main idea of this paper. On the left, we extend the plots from Figure 4 in Zhang et al. [65] by adding the MBBC variant and draw the BCO(0) and AIME separately. The purple region between the oracle version and the expert is the DKB while the orange region between the algorithm and its oracle version represents the EKB. On the right, we present the solutions proposed in this paper to solve the two barriers. The blue parts represent the original version of the algorithms that suffer from the knowledge barriers. Orange parts demonstrate the solution for EKB, where the agent is allowed to interact with the environment and use $D_{\mathrm{online}}$ together with $D_{\mathrm{body}}$ to update the world model. Purple parts show the solution for DKB, where a VIPER model is pretrained and used to label the online dataset and then used as a RL signal for policy learning.
@@
To better showcase the two barriers, in Figure 1 we extended Figure 4 from Zhang et al. [65] by adding a model-based BC (MBBC) variant where the policy is learned by running behaviour cloning on the learned latent space of the world model. BCO(0) [57] and AIME both pretrain a model from a large embodiment dataset and use that to infer the actions for the observation-only demonstrations. The oracle versions remove the need to infer the missing actions, thus removing the EKB. As we can see from the figure, the two algorithms are always upper-bounded by the corresponding oracle version, and the difference between them represents the EKB. On the other hand, even if given the true actions of the expert, imitation performance may still be impacted by a limited number of demonstrations providing insufficient coverage of the state space. Thus, the difference between the oracle version and the expert performance represents the DKB.
@@
In this paper, we study how to resolve these barriers to improve the performance of ILfO approaches, in particular of AIME proposed by Zhang et al. [65]. For the EKB, we extend the setting from offline to online by allowing the agent to further interact with the environment to gather more data to train the world model. While for the DKB, we introduce a surrogate reward function to allow the policy to essentially train on more data. We demonstrate that the proposed modifications significantly improve the performance on $9$ tasks in DeepMind Control Suite (DMC) [61] and $6$ tasks in Meta-World [64].
@@ 2 Preliminary
We mostly follow the problem setup as described in Zhang et al. [65]. We consider a POMDP problem defined by the tuple $\{S,A,T,R,O,\Omega\}$, where $S$ is the state space, $A$ is the action space, $T:S\times A\rightarrow S$ is the dynamic function, $R:S\rightarrow\mathbb{R}$ is the reward function, $O$ is the observation space, which is image in this paper, and $\Omega:S\rightarrow O$ is the emission function. The goal is to find a policy $\pi:S\rightarrow A$ which maximises the accumulated reward, i.e. $\sum_{t}r_{t}$.
We presume the existence of three datasets of the same embodiment available to our agent. The embodiment dataset $D_{\mathrm{body}}$ contains trajectories $\{o_{0},a_{0},o_{1},a_{1}\dots\}$ that represent past experiences of interacting with the environment. This dataset provides information about the embodiment for the algorithm to learn a world model. In addition, we also allow the agent to interact with the environment to collect new data in a replay buffer $D_{\mathrm{online}}$. Note that, although the simulator will give us the reward information, the agent is not allowed to use them, and we only use the reward for evaluation purpose. The demonstration dataset $D_{\mathrm{demo}}$ contains a few expert trajectories $\{o_{0},o_{1},o_{2}\dots\}$ of the embodiment solving a certain task defined by $R_{\mathrm{demo}}$. The crucial difference between this dataset and the other two datasets is that the actions are not provided anymore since they are not observable from a third-person perspective. The goal of our agent is to learn a policy $\pi$ from $D_{\mathrm{demo}}$ which can solve the task defined by $R_{\mathrm{demo}}$ as well as the expert who generated $D_{\mathrm{demo}}$.
@@ 2.1 World Models
A World Model [18] is a generative model which models a probability distribution over sequences of observations, i.e. $p(o_{1:T})$. The model can be either unconditioned or conditioned on other factors, such as previous observations or actions. When the actions taken are known, they can be considered as the condition, i.e.  $p(o_{1:T}|a_{0:T-1})$, and the model is called embodied [65].
In this paper, we consider variational latent world models where the observation is governed by a Markovian hidden state. In the literature, this type of model is also referred to as a State-Space Model (SSM) [30, 20, 19, 5, 33]. Such a variational latent world model involves four components, namely
@@
The objective function is composed of two terms: the first term $J^{\mathrm{rec}}$ is the likelihood of the observation under the inferred state, which is usually called the reconstruction loss; while the second term $J^{\mathrm{KL}}$ is the KL divergence between the posterior and the prior distributions of the latent state. To compute the objective function, we use the re-parameterisation trick [32, 50] to autoregressively sample the inferred states from the observation and action sequence.
@@
2.2 AIME
AIME is a recently proposed algorithm that uses a pretrained world model to solve ILfO in an offline setting. Specifically, it uses the pretrained world model as an implicit inference model by solving for the best action sequence that makes the demonstration most likely under the trained world model. The imitation can be done jointly with the action inference using amortised inference and the re-parameterisation trick by solving the following optimisation problem
@@ 3 Methodology
In this section we will make thorough analysis separately on the EKB and DKB especially upon the AIME algorithm as the case study. Based on the analysis we introduce the solution for each knowledge barriers and combine them as AIME-v2. We depict the general framework of the solutions on Figure 1 and show the pseudocode of AIME-v2 in Algorithm 1 in Appendix A.
3.1 Resolving the EKB
@@
The most natural way to solve the EKB is to allow the agent to further interact with the environment. By allowing the agent to interact with the environment, the new experiences can minimise the error in the pretrained model in proximity of the policy $\pi_{\psi}$ and gain more embodiment knowledge relevant for the task at hand. Torabi et al. [57] proposed a modified version of BCO(0) called BCO($\alpha$) which introduced such an interaction phase. However, from their and our empirical results, it did not resolve the EKB since there remains a gap between BCO($\alpha$) and the BC oracle when the environment is complex. In fact, as we will show in the following, the basic idea of adding online interactions is not straightforward to successfully implement in practice.
@@
As shown in recent works in Offline RL, continuing training an actor-critic from the offline phase in the online phase requires certain measures to combat the shift of objective [34, 4, 42]. A similar story also applies when extending AIME from purely offline to online. The most dominant problem we found is overfitting to the newly collected dataset.
@@
As the training progresses iteratively between data collection, model training and policy training, in the early phase of training, there are only a few new trajectories available for training the model. Because the world model is highly expressive, it may overly favor similar trajectories in the new data, leading to a high ELBO. Normally, this may not be a big problem since, eventually, more and more data will be collected to combat this overfitting. But since AIME also depends on the ELBO to train the policy, it quickly causes the policy training to diverge.
@@
In order to address the overfitting issue, we need a regulariser for model learning. Instead of designing ad-hoc methods to regularise the model in the parameter space, we adopt a data-driven approach. From model’s perspective, the overfitting is caused by a sudden shift of the training data from a large and diverse pretraining dataset to a small and narrow replay buffer. So one way to make the shift not as sudden is to append the pretraining dataset to the replay buffer, so that the distribution of the training data will change smoothly. However, this causes data efficiency problems since the replay buffer is relatively small compared to the pretraining dataset so that uniformly sampling from them together limits utilisation rate of the new data. Instead, we consider sampling separately from both datasets. Concretely, we modify the objective in Equation 2 to
@@
The amount of data we sample from the pretraining dataset is controlled by a hyper-parameter $\alpha$, which represents how much regularisation we put upon the model. In this paper, we mainly consider setting $\alpha=0.5$, so that we sample the data evenly from both datasets. We justify our choice in the ablation section with Figure 4(b).
@@
This finding contradicts Rajeswar et al. [48] and Hansen et al. [23], where the pretrained world models do not need such a data-driven regulariser. We conjecture that unlike AIME they mainly use their world models purely as generative models to predict states and rewards given action sequences, which is only indirectly influenced by overfitting the ELBO.
3.2 Resolving the DKB

### v2

# Overcoming Knowledge Barriers: Online Imitation Learning from Visual Observation with Pretrained World Models

@@ Abstract
Pretraining and finetuning models has become increasingly popular in decision-making. But there are still serious impediments in Imitation Learning from Observation (ILfO) with pretrained models. This study identifies two primary obstacles: the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB). The EKB emerges due to the pretrained models’ limitations in handling novel observations, which leads to inaccurate action inference. Conversely, the DKB stems from the reliance on limited demonstration datasets, restricting the model’s adaptability across diverse scenarios. We propose separate solutions to overcome each barrier and apply them to Action Inference by Maximising Evidence (AIME), a state-of-the-art algorithm. This new algorithm, AIME-NoB, integrates online interactions and a data-driven regulariser to mitigate the EKB. Additionally, it uses a surrogate reward function to broaden the policy’s supported states, addressing the DKB. Our experiments on vision-based control tasks from the DeepMind Control Suite and MetaWorld benchmarks show that AIME-NoB significantly improves sample efficiency and converged performance, presenting a robust framework for overcoming the challenges in ILfO with pretrained models. Code available at https://github.com/IcarusWizard/AIME-NoB.
@@ 1 Introduction
We have been going through a paradigm shift from learning from scratch to pretraining and finetuning, in particular in Computer Vision (CV) (He et al., 2016; Radford et al., 2021; He et al., 2022) and Natural Language Processing (NLP) (Devlin et al., 2019; Radford et al., ; Ouyang et al., 2022; Touvron et al., 2023a; b) fields due to the increasing availability of foundation models (Bommasani et al., 2021) and ever-growing datasets. However, it is still unclear how to adapt this new paradigm into decision-making, in particular what type of models we need to pretrain and how these models can be adapted to solve downstream tasks. Recent work (Zhang et al., 2023; DeMoss et al., 2023; Sekar et al., 2020; Rajeswar et al., 2023; Hansen et al., 2023a) showed that pretrained latent space world models enable successful and efficient transfer to new tasks with either reinforcement learning (Sekar et al., 2020; Rajeswar et al., 2023; Hansen et al., 2023a) or Imitation Learning from Observation (ILfO) (Zhang et al., 2023; DeMoss et al., 2023). ILfO (Torabi et al., 2018; 2019; Baker et al., 2022; Zhang et al., 2023; DeMoss et al., 2023; Liu et al., 2022a), especially from videos (Baker et al., 2022; Zhang et al., 2023; Liu et al., 2022a; DeMoss et al., 2023), is a more promising approach in this new paradigm since it does not require a handcrafted reward function which is hard to define for many real-world tasks.
@@
But there are challenges when using pretrained models in ILfO. To quantify these, we introduce two new barriers, which we call the Embodiment Knowledge Barrier (EKB) and the Demonstration Knowledge Barrier (DKB). The EKB describes the limitation of a pretrained model when confronted with novel observations and actions beyond its training experience. The DKB describes the generalisation from a limited number of expert demonstrations in imitation learning (Ho & Ermon, 2016). State-of-the-art approaches such as BCO(0) (Torabi et al., 2018) and AIME (Zhang et al., 2023) typically suffer from these two knowledge barriers. First, these algorithms depend on the pretrained model to infer missing actions from observation sequences. Thus, when the model has not seen a specific observation before, it may not know enough about the embodiment to infer the correct action. Second, if the policy optimisation is only guided by limited demonstrations, it can lead to a policy that generalises poorly, working well in some states but not in others.
@@
Figure 1: Main idea of this paper. On the left, we plot the performance of BCO(0) and AIME together with their oracle versions, which remove the EKB, w.r.t. different number of demonstrations on walker-run task. For each setting with the same number of the demonstrations, i.e. each column, the value difference between the oracle version and the expert is the Demonstration Knowledge Barrier (DKB) while the value difference between the algorithm and its oracle version represents the Embodiment Knowledge Barrier (EKB). On the right, we present the solutions proposed in this paper to overcome the two barriers. The blue parts represent the original version of the algorithms that suffer from the knowledge barriers. Orange parts demonstrate the solution for EKB, where the agent is allowed to interact with the environment and use $D_{\mathrm{online}}$ together with $D_{\mathrm{body}}$ to update the world model. Purple parts show the solution for DKB, where a surrogate reward model is derived from $D_{\mathrm{demo}}$ and used to label $D_{\mathrm{online}}$ and then used as an RL signal for policy learning.
@@
To better showcase the two barriers, in Figure 1, we evaluate both AIME (Zhang et al., 2023) and BCO(0) (Torabi et al., 2018) and their oracle versions w.r.t. different number of demonstrations on walker-run task. Both algorithms pretrain a model from a large embodiment dataset and use that to infer the actions for the observation-only demonstrations. The oracle versions remove the need to infer the missing actions by providing the algorithm with the true actions, thus removing the EKB. As we can see from the figure, for each setting with the same number of demonstrations, the two algorithms are always upper-bounded by the corresponding oracle version, and the difference between them represents the EKB. On the other hand, even if given the true actions of the expert, imitation performance may still be impacted by a limited number of demonstrations providing insufficient coverage of the state space. Thus, the difference between the oracle version and the expert performance represents the DKB.
@@
In this paper, we study how to overcome these barriers to improve the performance of ILfO approaches with pretrained models, in particular of AIME. For the EKB, we extend the setting from offline to online by allowing the agent to further interact with the environment to gather more data to train the world model. While for the DKB, we introduce a surrogate reward function to allow the policy to essentially train on more data. We demonstrate that the proposed modifications effectively overcome the two barriers and significantly improve the performance on nine tasks in DeepMind Control Suite (DMC) (Tunyasuvunakool et al., 2020) and six tasks in MetaWorld (Yu et al., 2021).
We summarise our contributions as follows:
  • We identify and thoroughly analyse the two knowledge barriers, namely EKB and DKB, in the current pretrained-model-based ILfO methods.
  • We propose AIME-NoB as an extension of the state-of-the art AIME algorithm by resolving the two knowledge barriers. Specifically, AIME-NoB uses online interaction with a data-driven regulariser to overcome the EKB and learn a surrogate reward function enlarging state coverage to overcome the DKB.
  • We evaluate AIME-NoB on 15 tasks from two vision-based benchmarks and the results demonstrate AIME-NoB significantly outperforms previous state-of-the-art methods both in terms of final performance and sample efficiency. We also conduct thorough ablation studies to show how the EKB and the DKB are overcome by the proposed modifications and how different design choices influence the performance.
@@ 2 Preliminary
We mostly follow the problem setup as described in Zhang et al. (2023). We consider a POMDP problem defined by the tuple $\{S,A,T,R,O,\Omega\}$, where $S$ is the state space, $A$ is the action space, $T:S\times A\rightarrow S$ is the dynamic function, $R:S\rightarrow\mathbb{R}$ is the reward function, $O$ is the observation space, and $\Omega:S\rightarrow O$ is the emission function. The goal is to find a policy $\pi:S\rightarrow A$ which maximises the expected accumulated reward, or return, i.e. $\mathcal{R}(\pi)=\mathbb{E}_{a\sim\pi}[\sum_{t=1}^{T}r_{t}]$. Since in this work we focus on imitation learning, this oracle reward is not available to the agent. We mainly use this reward to quantify the performance of the learnt policies.
We presume the existence of three datasets of the same embodiment available to our agent. The embodiment dataset $D_{\mathrm{body}}$ contains trajectories $\{o_{0},a_{0},o_{1},a_{1}\dots\}$ that represent past experiences of interacting with the environment. This dataset provides information about the embodiment for the algorithm to learn a world model. In addition, we also allow the agent to interact with the environment to collect new data in a replay buffer $D_{\mathrm{online}}$. Note that, although the simulator will give us the reward information, the agent is not allowed to use them, and we only use the reward for evaluation purposes. The demonstration dataset $D_{\mathrm{demo}}$ contains a few expert trajectories $\{o_{0},o_{1},o_{2}\dots\}$ of the embodiment solving a certain task defined by $R_{\mathrm{demo}}$. The crucial difference between this dataset and the other two datasets is that the actions are not provided anymore since they are not observable from a third-person perspective. The goal of our agent is to learn a policy $\pi$ from $D_{\mathrm{demo}}$ which can solve the task defined by $R_{\mathrm{demo}}$ as well as the expert $\pi_{\mathrm{demo}}$ who generated $D_{\mathrm{demo}}$.
@@ 2.1 World Models
A World Model (Ha & Schmidhuber, 2018) is a generative model which models a probability distribution over sequences of observations, i.e. $p(o_{1:T})$. The model can be either unconditioned or conditioned on other factors, such as previous observations or actions. When the actions taken are known, they can be considered as the condition, i.e. $p(o_{1:T}|a_{0:T-1})$, and the model is called embodied (Zhang et al., 2023). In this paper, we consider variational latent world models where the observation is governed by a Markovian hidden state. This type of model is also referred to as a State-Space Model (SSM) (Karl et al., 2017; Hafner et al., 2019b; a; Becker-Ehmck et al., 2019; Klushyn et al., 2021). Such a variational latent world model involves four components, namely
@@
The objective function is composed of two terms: the first term $J^{\mathrm{rec}}$ is the likelihood of the observation under the inferred state, which is usually called the reconstruction loss; while the second term $J^{\mathrm{KL}}$ is the KL divergence between the posterior and the prior distributions of the latent state. To compute the objective function, we use the re-parameterisation trick (Kingma & Welling, 2022; Rezende et al., 2014) to autoregressively sample the inferred states from the observation and action sequence.
@@
2.2 Action Inference by Maximising Evidence (AIME)
AIME is a state-of-the-art algorithm that uses a pretrained world model to solve ILfO in an offline setting. Specifically, it uses the pretrained world model as an implicit inference model by solving for the best action sequence that makes the demonstration most likely under the trained world model. The imitation can be done jointly with the action inference using amortised inference and the re-parameterisation trick by solving the following optimisation problem
@@ 3 Methodology
In this section we will analyse the EKB and DKB for AIME. Based on the analysis we introduce a solution for each knowledge barrier and combine them into AIME-NoB, where NoB stands for No Barriers. The general framework of the solutions is shown in Figure 1 and the pseudocode of AIME-NoB is provided in Algorithm 1. Before diving into the analysis, we first formally define EKB and DKB with
EKB $=\mathcal{R}(\pi_{\omega^{*}})-\mathcal{R}(\pi_{\psi^{*}}),$ (4)
DKB $=\mathcal{R}(\pi_{\mathrm{demo}})-\mathcal{R}(\pi_{\omega^{*}}).$ (5)
Let $\mathcal{D}{}\mathop{\mathrel{:}=}{}(D_{\mathrm{demo}},D_{\mathrm{body}},D_{\mathrm{online}})$, then
  • $\omega^{*}=\arg\!\max_{\omega}J_{\mathrm{policy}}(\hat{p}_{\pi_{\mathrm{demo}}}(a_{t}|o_{1:T}),\pi_{\omega}(a_{t}|o_{\leq t}),\mathcal{D})$ represents the optimal policy parameters for maximising $J_{\mathrm{policy}}$ with the oracle.
  • $\psi^{*}=\arg\!\max_{\psi}J_{\mathrm{policy}}(q_{\phi^{*}}(a_{t}|o_{1:T}),\pi_{\psi}(a_{t}|o_{\leq t}),\mathcal{D})$ represents the optimal policy parameters for maximising $J_{\mathrm{policy}}$ with the learned model.
  • $J_{\mathrm{policy}}(q(a_{t}|o_{1:T}),\pi(a_{t}|o_{\leq t}),\mathcal{D})$ is the learning objective for the policy depending on an action-inference model $q(a_{t}|o_{1:T})$. For behaviour cloning based methods like BCO and AIME, it is essentially equivalent to $-\sum_{o_{1:T}\in D_{\mathrm{demo}}}\sum_{t}D_{\text{KL}}(q(a_{t}|o_{1:T})\,|\,\pi(a_{t}|o_{\leq t}))$.
  • $\phi^{*}=\arg\!\max_{\phi}J_{\mathrm{model}}(D_{\mathrm{body}},D_{\mathrm{online}},\phi)$ is the optimal parameter for maximising the model-learning objective $J_{\mathrm{model}}$.
  • $\hat{p}_{\pi_{\mathrm{demo}}}(a_{t}|o_{1:T})$ is the ground truth of empirical distribution of the demonstration data that serves as an oracle.
Based on the definitions, we can clearly see that EKB is caused by the imperfect inference of the unknown actions of the observation-only demonstrations, which can be characterized by $D_{\text{KL}}(\hat{p}_{\pi_{\mathrm{demo}}}(a_{1:T-1}|o_{1:T})\,|\,q_{\phi^{*}}(a_{1:T-1}|o_{1:T}))$; while DKB is caused by the limited learning signal that $J_{\mathrm{policy}}$ offers as it is only defined on the data points in the demonstration dataset. In Appendix J, we propose some upper bounds for both EKB and DKB based on these insights. In the rest of this section, we will focus on practical solutions for both of the barriers.
3.1 Overcoming the EKB
@@
To reduce the EKB, we need to bring the learned inference model closer to the oracle, i.e. to minimise $\sum_{o_{1:T}\in D_{\text{demo}}}D_{\text{KL}}(\hat{p}_{\pi_{\mathrm{demo}}}(a_{1:T-1}|o_{1:T})\,|\,q_{\phi^{*}}(a_{1:T-1}|o_{1:T}))$. The most natural way to do so is to allow the agent to further interact with the environment, similar to how we humans practice for a novel skill. New experiences can minimise the error in the pretrained model near the policy $\pi_{\psi}$ and enhance task-specific embodiment knowledge. Torabi et al. (2018) proposed a modified version of BCO(0) called BCO($\alpha$) introducing such an interaction phase. However, empirical results show it did not overcome the EKB as a gap remains with the BC oracle when the environment is complex. In fact, as we will show in the following, the idea of adding online interactions is not straightforward to successfully implement in practice.
@@
As shown in recent works in Offline RL, continuing training of an actor-critic from the offline phase in the online phase requires certain measures to combat the shift of objective (Lee et al., 2022; Ball et al., 2023; Nakamoto et al., 2023). A similar story also applies when extending AIME from purely offline to online. The most dominant problem we found is overfitting to the newly collected dataset.
@@
As the training progresses alternating between data collection, model training and policy training, in the early phase of training there are only very few new trajectories available for training the model. Because the world model is highly expressive, it may overly favour similar trajectories, especially the action sequence, in the new data, leading to a high ELBO. Normally, this may not be a big problem since, eventually, more and more data will be collected to combat this overfitting. But since AIME also depends on the ELBO to train the policy, it quickly causes the policy training to diverge. That is to say, when the model is extensively trained on a small amount of data, it not only maximises the conditional likelihood $\log(o_{1:T}|a_{0:T-1})$ but also maximises the marginal likelihood $\log(\cdot|a_{0:T-1})$, which diverges the likelihood-based action inference process.
@@
In order to address the overfitting issue, we need a regulariser for model learning. Instead of designing ad-hoc methods to regularise the model in the parameter space, we adopt a data-driven approach. From the model’s perspective, the overfitting is caused by a sudden shift of the training data from a large and diverse pretraining dataset to a small and narrow replay buffer. To avoid this sudden shift, the most straightforward method is to just append new data to a replay buffer that is pre-filled with the pretraining dataset. However, this causes data efficiency problems since the newly collected data is relatively little compared to the big pretraining dataset. Uniformly sampling from the joint replay buffer hence overly limits usage of the new data. Instead, we suggest sampling separately from both datasets. We modify the model-learning objective $J_{\mathrm{model}}$ in Equation 2 to
@@
The amount of data we sample from the pretraining dataset is controlled by a hyper-parameter $\alpha$, which represents how much regularisation we put upon the model. Here we mainly consider setting $\alpha=0.5$, so that we sample the data evenly from both datasets.
@@
This finding contradicts Rajeswar et al. (2023) and Hansen et al. (2023a), where the pretrained world models do not need such a data-driven regulariser. We conjecture that unlike AIME, these approaches mainly use their world models purely as generative models to predict states and rewards given action sequences, which is only indirectly influenced by overfitting the ELBO.
Algorithm 1 AIME-NoB
1:  Input: Embodiment dataset $D_{\mathrm{body}}$, Demonstration dataset $D_{\mathrm{demo}}$, Pretrained world model parameters $\phi,\theta$, Surrogate reward model $R_{\nu}$, Regulariser ratio $\alpha$, Value gradient weight $\beta$, Batch size $\mathrm{B}$
2:  Initialise policy and critic parameters $\psi,\xi$ randomly.
3:  for $i=1$ to policy pretraining iterations do
4:     Draw a batch of demonstrations $o_{1:T}\sim D_{\mathrm{demo}}$.
5:     Update policy parameters $\psi$ with Equation 3.
6:  end for
7:  Initialize $D_{\mathrm{online}}\rightarrow\emptyset$.
8:  for $i=1$ to Environment Interaction budget do
9:     Collect a new episode $\{o_{1:T},a_{1:T}\}$ with the current policy $\pi_{\psi}$
10:     Update the surrogate reward model $R_{\nu}$ if needed, e.g. for AIL.
11:     Estimate reward using surrogate reward model $r_{1:T}^{\mathrm{sur}}=R_{\nu}(o_{1:T})$
12:     Append $\{o_{1:T},a_{1:T},r_{1:T}^{\mathrm{sur}}\}$ to $D_{\mathrm{online}}$
13:     # Update world model
14:     Draw $\alpha\cdot\mathrm{B}$ samples $b_{\mathrm{body}}\sim D_{\mathrm{body}}$
15:     Draw $(1-\alpha)\cdot\mathrm{B}$ samples $b_{\mathrm{online}}\sim D_{\mathrm{online}}$
16:     Define combined batch $b=b_{\mathrm{body}}\cup b_{\mathrm{online}}$
17:     Finetune model with batch $b$ using Equation 6.
18:     # Update policy
19:     Sample a batch from $D_{\mathrm{demo}}$
20:     Update policy parameters $\psi$ with Equation 9.
21:     Update value function parameters $\xi$ with Equation 8.
22:  end for
3.2 Overcoming the DKB